OpenStudy (mathmath333):

what power of 8 exactly divides 25! ?

ganeshie8 (ganeshie8):

hint : 8 = 2^3

OpenStudy (anonymous):

ok if 8 is prime, answer would have been 25/8=3 but since it is not do like this.....

OpenStudy (anonymous):

ok... 8=2*2*2. for divisibility by 8 we need three 2's.We add one to the power of 8 which divide 25!.TO count how we get 3 two's we do the following..... All even no.s will give 1 two's at least 25/2=12 also all no.s in 25! is divisible by 2^2 will give us additional two's 25/2^2=6 further all no. in 25! is divisible by2^3------->3 and with 2^4------->1 so total number is 22,the answer.

OpenStudy (anonymous):

Hope it's clear!! :O

OpenStudy (mathmath333):

does it has any other method

OpenStudy (anonymous):

well idk, it's only the easiest method i know :D

OpenStudy (anonymous):

since 25! has 22 two's it will be divided by 8 seven times. :)

OpenStudy (anonymous):

Answer is 8^7......what's the answer?

OpenStudy (mathmath333):

(25/2)+(25/4)+(25/8)+(25/16) its easiest

OpenStudy (anonymous):

oh.!! well then that's nice. Good for you.Thanks for the advice.

OpenStudy (anonymous):

what i said is the same.....you just showed it in one go :)